A canon ball is shot horizontally from a point 24 meters above the ground and strikes the ground after traveling a distance of 55 meters. With what speed was it shot?
1
Expert's answer
2012-09-25T11:16:01-0400
h = g * t^2 / 2, s = V * t
t = sqrt( 2 * h / g )
V = s / t = s / sqrt( 2 *
h / g )
V = 55 / sqrt( 2 * 24 / 9.81 ) = 24.86 m/s
Answer: 24.86 m/s.